# A comparative high-resolution spectroscopic analysis of in situ and accreted globular clusters

### v1

@@ Acknowledgements.
Based on observations collected at the ESO-VLT under the programs 069.D-0227 (P.I. P. François), 073.D-0211 (P.I. E. Carretta), 083.D-0208 (P.I. E. Carretta), 097.D-0175 (PI: B. Barbuy), 091.D-0535 (PI: C. Moni Bidin), 0101.D-0109 (P.I. A. Marino), 193.D-0232 (P.I. F. R. Ferraro), and 197.B-1074 (P.I. G. F. Gilmore). This research is funded by the project LEGO – Reconstructing the building blocks of the Galaxy by chemical tagging (P.I. A. Mucciarelli), granted by the Italian MUR through contract PRIN 2022LLP8TK_001. MB, EC, AM and DM acknowledge the support to this study by the PRIN INAF 2023 grant ObFu CHAM - Chemodynamics of the Accreted Halo of the Milky Way (P.I. M. Bellazzini). DM acknowledges financial support from PRIN-MIUR-22 “CHRONOS: adjusting the clock(s) to unveil the CHRONO-chemo-dynamical Structure of the Galaxy” (PI: S. Cassisi) granted by the European Union - Next Generation EU. DM and MB acknowledge the support to activities related to the ESA/Gaia mission by the Italian Space Agency (ASI) through contract 2018-24-HH.0 and its addendum 2018-24-HH.1-2022 to the National Institute for Astrophysics (INAF). We thank the anonymous referee for the helpful comments that improved the quality of the paper. This work has made use of data from the European Space Agency (ESA) mission Gaia https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This work made use of SDSS-IV data. Funding for the Sloan Digital Sky Survey IV has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions. SDSS-IV acknowledges support and resources from the Center for High Performance Computing at the University of Utah. The SDSS website is www.sdss4.org. SDSS-IV is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics — Harvard & Smithsonian, the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatário Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University.References
@@
- François et al. (2024) François, P., Cescutti, G., Bonifacio, P., et al. 2024, arXiv e-prints, arXiv:2404.08418
@@
- Monty et al. (2024) Monty, S., Belokurov, V., Sanders, J. L., et al. 2024, arXiv e-prints, arXiv:2405.08963
@@ A.1 GSE globular clusters
To validate our findings, we undertake a comparative analysis with two distinct independent studies conducted on two of the six target GCs. Specifically, we compare our results to those obtained by the seminal works by Carretta et al. (2009, 2010, 2013) and to those provided by the latest data release of the APOGEE survey (Abdurro’uf et al. 2022). In particular, we use the APOGEE value-added catalogue of Galactic GC stars (Schiavon et al. 2024), limiting our comparison only to well-measured stars. To do so, we remove all the stars with the following flags333Flags definition available at https://www.sdss4.org/dr17/irspec/apogee-bitmasks/. in the quality parameters:
@@
The choice of using these two dataset was guided by their execution of identical chemical analyses on two out of the four target GCs, thereby ensuring consistency in their outcomes. We limit our comparison with Carretta et al. (2009, 2010, 2013) and APOGEE only to the chemical elements identified as sensitive to the origin of GCs available in these dataset, that are Mg, Si and Ca. As shown in Figure 11, the results from this work are in agreement within the uncertainties with those provided by Carretta et al. (2009, 2010, 2013) and APOGEE, consistently yielding an enhancement of - in the -elements (Mg, Si, and Ca) of NGC 6218 with respect to NGC 362.
@@ A.2 Comparison with V13
A comparison with the results of V13, who analyzed the same Rup106 dataset, further confirms the peculiarity of this system. Indeed, their findings indicate that, within this cluster, the observed spread in all chemical elements is consistent with the uncertainty on the measurement, suggesting Rup106 to be a single population GC, as also confirmed photometrically in more recent works (Dotter et al. 2018; Lagioia et al. 2024). Moreover, their analysis reveals solar-like abundances in the -elements, alongside a notable underabundance in the iron peak elements relative to iron, ranging from up to dex. Firstly, when comparing the atmospheric parameters of Rup106 targets, we find a significant difference in the adopted scales, with an average discrepancy (this work – V13) of K ( K). The average difference for the surface gravity is () and that for is km/s ( km/s). These discrepancies arise from the fact that V13 constrained and log with a metodology similar to ours, deriving them from the photometry, but exploiting a different color - relation. On the other hand, V13 assumed a derived from the relation by Marino et al. (2008), while we constrained it spectroscopically (see Section 3.1). To ensure a fair comparison between the results, we rescaled the abundance ratios reported in V13 on the Grevesse & Sauval (1998) solar composition. Firstly, we note an average offset of ( = ) in the star-to-star metallicity estimate of Rup106, which can be explained by the difference in the derived effective temperatures. By comparing these two independent results performed on the same dataset, we find agreement with V13 at the level for Mg, Ti, Sc, Cr, Mn, Zn, Y, La, and Eu, while we note differences up to for Si, Ca, V, Co, Ni, Cu, and Ba. We attribute these inconsistencies to the different scale adopted for surface gravities and, most importantly, effective temperatures. This comparison proves that diverse assumptions in the procedure to derive chemical abundances make the results of two independent approaches not directly comparable and that the only reliable chemical analysis is within an homogeneous procedure. In general, notwithstanding differences stemming from varied assumptions in the chemical analysis, we reaffirm the distinct chemical characteristics of Rup106 within the MW GC system, alongside the minimal dispersion in chemical abundances among its constituent stars.

### v2

@@ Acknowledgements.
Based on observations collected at the ESO-VLT under the programs 069.D-0227 (P.I. P. François), 073.D-0211 (P.I. E. Carretta), 083.D-0208 (P.I. E. Carretta), 091.D-0535 (PI: C. Moni Bidin), 097.D-0175 (PI: B. Barbuy), 0101.D-0109 (P.I. A. Marino), 193.D-0232 (P.I. F. R. Ferraro), and 197.B-1074 (P.I. G. F. Gilmore). This research is funded by the project LEGO – Reconstructing the building blocks of the Galaxy by chemical tagging (P.I. A. Mucciarelli), granted by the Italian MUR through contract PRIN 2022LLP8TK_001. MB, EC, AM and DM acknowledge the support to this study by the PRIN INAF 2023 grant ObFu CHAM - Chemodynamics of the Accreted Halo of the Milky Way (P.I. M. Bellazzini). DM acknowledges financial support from PRIN-MIUR-22 “CHRONOS: adjusting the clock(s) to unveil the CHRONO-chemo-dynamical Structure of the Galaxy” (PI: S. Cassisi) granted by the European Union - Next Generation EU. DM and MB acknowledge the support to activities related to the ESA/Gaia mission by the Italian Space Agency (ASI) through contract 2018-24-HH.0 and its addendum 2018-24-HH.1-2022 to the National Institute for Astrophysics (INAF). We thank the anonymous referee for the helpful comments that improved the quality of the paper. This work has made use of data from the European Space Agency (ESA) mission Gaia https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This work made use of SDSS-IV data. Funding for the Sloan Digital Sky Survey IV has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions. SDSS-IV acknowledges support and resources from the Center for High Performance Computing at the University of Utah. The SDSS website is www.sdss4.org. SDSS-IV is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics — Harvard & Smithsonian, the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatário Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University.References
@@
- François et al. (2024) François, P., Cescutti, G., Bonifacio, P., et al. 2024, A&A, 686, A295
@@
- Monty et al. (2024) Monty, S., Belokurov, V., Sanders, J. L., et al. 2024, MNRAS, 533, 2420
@@ A.1 GSE globular clusters
To validate our findings, we undertake a comparative analysis with two distinct independent studies conducted on two of the six target GCs. Specifically, we compare our results to those obtained by the seminal works by Carretta et al. (2009, 2010, 2013) and to those provided by the latest data release of the APOGEE survey (Abdurro’uf et al. 2022). In particular, we use the APOGEE value-added catalogue of Galactic GC stars (Schiavon et al. 2024), limiting our comparison only to well-measured stars. To do so, we remove all the stars with the following flags888Flags definition available at https://www.sdss4.org/dr17/irspec/apogee-bitmasks/. in the quality parameters:
@@
The choice of using these two dataset was guided by their execution of identical chemical analyses on two out of the four target GCs, thereby ensuring consistency in their outcomes. We limit our comparison with Carretta et al. (2009, 2010, 2013) and APOGEE only to the chemical elements identified as sensitive to the origin of GCs available in these dataset, that are Mg, Si and Ca. As shown in Fig. 11, the results from this work are in agreement within the uncertainties with those provided by Carretta et al. (2009, 2010, 2013) and APOGEE, consistently yielding an enhancement of - in the -elements (Mg, Si, and Ca) of NGC 6218 with respect to NGC 362.
@@ A.2 Comparison with V13
A comparison with the results of V13, who analyzed the same Rup106 dataset, further confirms the peculiarity of this system. Indeed, their findings indicate that, within this cluster, the observed spread in all chemical elements is consistent with the uncertainty on the measurement, suggesting Rup106 to be a single population GC, as also confirmed photometrically in more recent works (Dotter et al. 2018; Lagioia et al. 2024). Moreover, their analysis reveals solar-like abundances in the -elements, alongside a notable underabundance in the iron peak elements relative to iron, ranging from up to dex. Firstly, when comparing the atmospheric parameters of Rup106 targets, we find a significant difference in the adopted scales, with an average discrepancy (this work – V13) of K ( K). The average difference for the surface gravity is () and that for is km/s ( km/s). These discrepancies arise from the fact that V13 constrained and log with a metodology similar to ours, deriving them from the photometry, but exploiting a different colour - relation. On the other hand, V13 assumed a derived from the relation by Marino et al. (2008), while we constrained it spectroscopically (see Sect. 3.1). To ensure a fair comparison between the results, we rescaled the abundance ratios reported in V13 on the Grevesse & Sauval (1998) solar composition. Firstly, we note an average offset of ( = ) in the star-to-star metallicity estimate of Rup106, which can be explained by the difference in the derived effective temperatures. By comparing these two independent results performed on the same dataset, we find agreement with V13 at the level for Mg, Ti, ScII, Cr, Mn, Zn, YII, LaII, and EuII, while we note differences up to for Si, Ca, V, Co, Ni, Cu, and BaII. We attribute these inconsistencies to the different scale adopted for surface gravities and, most importantly, effective temperatures. This comparison proves that diverse assumptions in the procedure to derive chemical abundances make the results of two independent approaches not directly comparable and that the only reliable chemical analysis is within an homogeneous procedure. In general, notwithstanding differences stemming from varied assumptions in the chemical analysis, we reaffirm the distinct chemical characteristics of Rup106 within the MW GC system, alongside the minimal dispersion in chemical abundances among its constituent stars.